OpenStudy (perii224):

PLEASE HELP WILL MEDAL Determine if the sequence is arithmetic or geometric. Then find the next term in the sequence. Select all that apply. 1, -4, 16, -64, … arithmetic geometric 128 256 Find the common ratio, r, in the following geometric sequence. 2, 6, 18, 54, … 3 -3 4 9

OpenStudy (perii224):

@welshfella can you help ?

OpenStudy (welshfella):

for the first one try dividing the first term by the second then 3rd by 2nd etc and see if you get the same result first / 2nd = 1 / -4 = 1/4 2nd / 3rd = -4/16 = ?

OpenStudy (welshfella):

if these are all the same then we have a common ratio and the series is geometric.

OpenStudy (welshfella):

* that is -1/4 not 1/4

OpenStudy (welshfella):

-4/16 = what?

OpenStudy (perii224):

-0.25?

OpenStudy (welshfella):

yes or -1/4 and finally 16/-64 = -1/4 so its geometric

OpenStudy (perii224):

Oh okay, thank you! :)

OpenStudy (welshfella):

Oh - SIlli me!! we should be doing the opposite division - like -4 / 1 = -4 The common ratio is -4 not -1/4. - the common ration multiplies each term to get the next so the next term in series is -64 * -4 = 256

OpenStudy (welshfella):

for second question r = 6/2

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

check this is true by dividing the others 18/6 = 54/18 =

OpenStudy (welshfella):

the general form of a geometric series is a, ar , ar^2, ar^3 , ar^4..... a is the first term and r = common ratio

OpenStudy (welshfella):

gotta go right now

OpenStudy (perii224):

Thanks so much! :)